OpenStudy (anonymous):

g(x)= 4x^-5, h(x)= (x+2)/4 what is g(x)/h(x), please she work/explain g(x)= 4x^-5, h(x)= (x+2)/4 what is g(x)/h(x), please she work/explain @Mathematics

OpenStudy (anonymous):

i need it as if it still had X in the equation

OpenStudy (anonymous):

g(x)= 4x^-5, h(x)= (x+2)/4 try to divide the g/h

OpenStudy (anonymous):

well, 1st off, if you take the negative exponent and make it a positive by making it a detonator, does the 4 go with it?

OpenStudy (anonymous):

where does the 16 come from?

OpenStudy (anonymous):

the 4 in the denominator will be mult by the numerator 4

OpenStudy (anonymous):

ok, so if we do go with that, then the type of answer im looking for is something like 16(1/x^6+2)

OpenStudy (anonymous):

no, because what did you do to get rid of the negative?

OpenStudy (anonymous):

the x^-5 wil be mult by the denominator when you changed to positive sign

OpenStudy (anonymous):

i was taught that it moves to the detonator

OpenStudy (anonymous):

yes look at my solution now,,it goes to the denom and becomes positive on sign

OpenStudy (anonymous):

4x^-5, h(x)= (x+2)/4 =16/[x^5(x+2)]=16/[(x^6 +2x^5)] ans

OpenStudy (anonymous):

the 16 remains on the numerator and the x^5 now in the denom

OpenStudy (anonymous):

but theres only one x^-5 and one x as demonators, so how are there a total of 11 in the end?

OpenStudy (anonymous):

note on the sign of16(x^-5)/(x+2)

OpenStudy (anonymous):

16(x^-5)/(x+2) =16/[x^+5)(x+2)]

OpenStudy (anonymous):

ohhh, you took that as a minus!!! its a negative

OpenStudy (anonymous):

now you can mult the denominator

OpenStudy (anonymous):

yes im sory i didnot noticed the negative sign before, now its like this 16(x^-5)/(x+2) =16/[x^+5)(x+2)] =16/[(x^6 +2x^5)] ans

OpenStudy (anonymous):

did you get it jiwani?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

tell me if you get confused for i will explain it

OpenStudy (anonymous):

aroub may explain it better than me,....lol

OpenStudy (aroub):

haaha!! no its all yours man!! =D

OpenStudy (anonymous):

i have a test over this stuff in a few hours so i dont know how much of it my teacher will explain to me, if any at all. but idk, im just gonna ask her and see if she can explain it

OpenStudy (anonymous):

ok thnx aroub, i gotta go now,,,lol

OpenStudy (anonymous):

ok aroub its your turn

OpenStudy (anonymous):

hm jiwani believed me my solution there is correct,,try to se what happened to x^-5 went down to the denom. and changed to x^5